Assuming the density of nitrogen at s.t.p to be 1.251 kgm^-3, find the root mean square velocity of nitrogen molecules at 127 oC.
1
Expert's answer
2021-02-09T13:33:30-0500

"T = 127 + 273 = 400 \; K \\

M(N_2) = 28\; g/mol = 28 \times 10^{-3} \;kg/mol \\

V_{rms} = \sqrt{\frac{3RT}{M}} \\

= \sqrt{\frac{3 \times 8.314 \times 400}{28 \times 10^{-3}}} \\

= \sqrt{356314.28} \\

= 596.9 \;m/s"

Answer: 596.9 m/s